OpenStudy (anonymous):

Mike started a savings account by depositing $9. Each month he deposits more money than the month before. At the end of 41 months, he has saved $9,389.00. How much more does he deposit each month?

OpenStudy (amistre64):

'more money' is vague. but i spose we could work a general idea from it

OpenStudy (amistre64):

nothing is mentioned about interest so i assume this is under his mattress?

OpenStudy (amistre64):

its most likely an arithmetic progression summed up over 41 months

OpenStudy (anonymous):

I understand arithmetics, but I don't get this specific question. I don't think that interest is part of the question.

OpenStudy (amistre64):

i dont spose you recall the formula for the sum of an arithmetic sequence?

OpenStudy (anonymous):

Yeah. One sec.

OpenStudy (anonymous):

There are two. One is Sn=(n/2)(a_1 + a_n). That one is simpler.

OpenStudy (anonymous):

The other one is Sn=(n/2)[2a_1 + (n-1)d].

OpenStudy (amistre64):

good, all you need now is to determine what a_41 would be ... a_n = a_1 + d(n-1) which when subbed in gives the 2nd one you posted Sn=(n/2)(a_1 + a_n) Sn=(n/2)(a_1 + (a_1 + d(n-1))) Sn=(n/2)(a_1 + a_1 + d(n-1)) Sn=(n/2)( 2a_1 + d(n-1))

OpenStudy (amistre64):

we know a_1 and n; and S_41 ... so solve for d

OpenStudy (anonymous):

But, it says that he added "more money than the month before." Does this mean that there is exponential growth?

OpenStudy (amistre64):

let call Sn, K to avoid confusing n parts ... K = (n/2) (2 a1 + d(n-1)) 2K/n = 2 a1 + d(n-1) 2K/n - 2 a1 = d(n-1) 2[ K/n - a1] = d(n-1) 2[ K/n - a1]/(n-1) = d

OpenStudy (amistre64):

exponential is a possibility but that was why i said it was vague to start with.

OpenStudy (anonymous):

Let's leave that out if we can because it did not tell me to do that.

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

2[ K/n - a1]/(n-1) = d? So, what goes in there? Let me re-check my lesson to see if there is a simpler way real quick.

OpenStudy (amistre64):

lol, n=41 is given, a1 =9 is given ... and K is used to avoid overuse of n and is just the end amount he saved up

OpenStudy (amistre64):

..which is given

OpenStudy (anonymous):

So, 2[ 9,389/n - 9]/(41-1) = d?

OpenStudy (amistre64):

yes, but you missed an 'n' :)

OpenStudy (amistre64):

2[ 9,389/41 - 9]/(41-1) 2[ 9,389/41 - 9]/(40) (9,389/41 - 9)/20 etc ...

OpenStudy (anonymous):

Why did the 40 go to a 20?

OpenStudy (amistre64):

2/40 = 1/20 of course :/

OpenStudy (anonymous):

I thought that it was being multiplied by 2.

OpenStudy (amistre64):

write it down on paper if you have to ... \[\large \frac{2(\frac{9389}{41}-9)}{41-1}\] \[\large \frac{\cancel2^1(\frac{9389}{41}-9)}{\cancel{40}_{20}}\]

OpenStudy (anonymous):

My options are: $11.00 $11.50 $12.00 $12.50

OpenStudy (amistre64):

yeah, we get one of those ...

OpenStudy (anonymous):

I got about 14...?

OpenStudy (anonymous):

Oh, no. It was my own stupid fault. I put the data in wrong. My bad.

OpenStudy (amistre64):

:) good luck and all. i gotta meet a guy about a thing

OpenStudy (anonymous):

Thanks for your help. Have a good one.